# Spatially parallel decoding for multi-qubit lattice surgery

### v1

@@ Abstract
Running quantum algorithms protected by quantum error correction requires a real time, classical decoder. To prevent the accumulation of a backlog, this decoder must process syndromes from the quantum device at a faster rate than they are generated. Most prior work on real time decoding has focused on an isolated logical qubit encoded in the surface code. However, for surface code, quantum programs of utility will require multi-qubit interactions performed via lattice surgery. A large merged patch can arise during lattice surgery – possibly as large as the entire device. This puts a significant strain on a real time decoder, which must decode errors on this merged patch and maintain the level of fault-tolerance that it achieves on isolated logical qubits.
@@ III Related work
The work in [4] is the only one prior to ours that targets lattice-surgery style fault-tolerant blocks. But its focus is different from ours. It prioritizes minimizing the latency of a software implementation and does not address “further hardware and systems considerations that are relevant”. We instead prioritize the constraints of hardware accelerators — which are likeliest to meet the throughput requirements. For example, they develop algorithms that take lattice surgery blocks as input and assigns each to a window, while we insist that windows have pre-assigned positions. As another contrast, they focus on showing the accuracy of the parallel scheme when the buffer width is at least $d$. Meanwhile, we study how small one can make the buffer width while still achieving high accuracy.
@@ VI-B Inner decoders
An inner decoder’s scalability bottleneck depends on its design. As examples, we look at two FPGA-based real time decoders that work for at least medium-sized codes and have different scalability challenges. Helios [28] is a distributed version of a Union-Find decoder that exploits parallel computing resources for speedup. It has been demonstrated on a $21\times 21$ patch. It operates with $O(d^{3})$ processing elements, each assigned to a node in the decoding graph and communicates via shared memory. Its average runtime per measurement round decreases with $d$, as long as $d$ is not large enough to require board to board communication between FPGAs (which is much slower than shared memory based communication). Aside from this limit, the main scalability challenge of Helios is the $O(d^{3})$ scaling of its computing resources.
@@
In the two examples, the maximum window size supported by the FPGA is constrained by the scalability of the inner decoder. Both decoders work for at least $21^{3}$ decoding graphs, though. For the spatially parallel windows, this translates to a commit region width that is at least $10$, assuming that the buffer is slightly wider than half of the commit region. Furthermore, the decoding graphs in [2, 28] are cubes, but for the spatially parallel windows, the number of measurement rounds that need to be processed at once does not depend on the full width of the window, but rather the width of the commit region. In the case of $w\approx d/2$ (so that a full window has width $2d$ while the commit region has width $d$), this means the volume of the decoding graph is not $(2d)^{3}=8d^{3}$ but $(2d)^{2}\times d=4d^{3}$. Thus, the inner decoders should be able to work on larger windows than the sizes of isolated patches that they are reported to support. We conclude that FPGA-based implementations of real time decoders connected into spatially parallel windows will be able to support short to medium term demonstrations of logical operations with $d$ up to $11$ or $\sim 15$. In the long term, ASIC-based implementations should be able to support larger code distances, because they can be optimized for higher performance and lower cost [2].

### v2

@@ Abstract
Running quantum algorithms protected by quantum error correction requires a real time, classical decoder. To prevent the accumulation of a backlog, this decoder must process syndromes from the quantum device at a faster rate than they are generated. Most prior work on real time decoding has focused on an isolated logical qubit encoded in the surface code. However, for surface code, quantum programs of utility will require multi-qubit interactions performed via lattice surgery. A large merged patch can arise during lattice surgery — possibly as large as the entire device. This puts a significant strain on a real time decoder, which must decode errors on this merged patch and maintain the level of fault-tolerance that it achieves on isolated logical qubits.
@@ III Related work
The work in [4] is the only one prior to ours that targets lattice-surgery style fault-tolerant blocks. Compared to it, we put more emphasis on the constraints of hardware accelerators — which are likeliest to meet the throughput requirements. For example, we force the windows to have pre-assigned positions and layers. Additionally, both papers include numerical results on how the buffer width affects decoding accuracy, but we provide more analysis and also study how the influence depends on the physical noise level.
@@ VI-B Inner decoders
An inner decoder’s scalability bottleneck depends on its design. As examples, we look at two FPGA-based real time decoders that work for at least medium-sized codes and have different scalability challenges. Helios [29] is a distributed version of a Union-Find decoder that exploits parallel computing resources for speedup. It has been demonstrated on a $21\times 21$ patch and operates with $O(d^{3})$ processing elements, each assigned to a node in the decoding graph and communicates via shared memory. Its average runtime per measurement round decreases with $d$, as long as $d$ is not large enough to require board to board communication between FPGAs (which is much slower than shared memory based communication). Aside from this limit, the main scalability challenge of Helios is the $O(d^{3})$ scaling of its computing resources. A recent update of Helios [20] improved its scalability to a $51\times 51$ patch at the cost of increased latency.
@@
The maximum window size supported by the FPGA is constrained by the scalability of the inner decoder. The decoding graphs in [2, 29] are cubes, but for the spatially parallel windows, the number of measurement rounds that need to be processed at once does not depend on the full width of the window, but rather the width of the commit region. This would decrease the volume of the decoding graph by roughly half, but in practice, applying sliding windows along the time dimension would increase the volume of the decoding graph by $2X$. Therefore we expect that volume of the decoding graph in a window would be similar to the ones used in [2, 29], which means an inner decoder for $d=51$ translates to a commit region width of $\sim 25$ for the spatially parallel windows, assuming that the buffer is slightly wider than half of the commit region. We conclude that FPGA-based implementations of real time decoders connected into spatially parallel windows will be able to support at least medium term demonstrations of logical operations. In the long term, ASIC-based implementations should be able to support larger code distances, because they can be optimized for higher performance and lower cost [2].